### 09 Apr 2026 v2

# Equivalent Circuit Modeling of Foil-Mediated Dissipative Coupling in Microwave Cavities with Enhanced Phase Response

Abstract
We formulate and validate an equivalent circuit model describing mutual resistive coupling between three microwave cavity resonators interconnected via thin metallic foils. Each cavity is represented as a lumped LCR circuit, while the foils act as a dissipative interface that mediates energy exchange via mutual resistance. This coupling mechanism produces interference effects and a controllable anti-resonance when the input resonators are amplitude- and phase-balanced, a behavior not achievable with standard microwave antenna probes. All three resonators operated in the TM010 mode, where two input resonators each excited the third via a thin copper foil. Analytical expressions are derived for the mutual resistance and coupling coefficient of these foils in this geometry. Under balanced conditions, a sharp anti-resonance emerges with a near order-of-magnitude enhanced phase sensitivity at the resonant frequency of the output cavity, consistent with model predictions. The experimentally extracted mutual coupling coefficients, $\Delta_{13}=(5.00\pm 0.01)\times 10^{-6}$ and $\Delta_{23}=(4.10\pm 0.01)\times 10^{-6}$, fall within the calculated range $\Delta_{n3}\approx(1\text{--}48)\times 10^{-6}$ derived from the foil’s electromagnetic properties, where the spread is dominated by the estimated foil thickness uncertainty of $(9\pm 1)\,\mu\mathrm{m}$. These results confirm that resistive coupling can occur across a number of skin depths of a metallic interface, providing a new means of engineering controlled interference in multi-resonator systems. The approach offers potential applications in precision microwave experiments, phase-sensitive detection, and tests of fundamental electromagnetic interactions.
I Introduction
Microwave cavity resonators are foundational components in precision measurement systems [18], frequency standards [21], and emerging quantum technologies [24, 12, 19]. Their ability to confine and sustain electromagnetic fields with extremely high quality factors makes them indispensable for experiments requiring long coherence times and minimal dissipation [26]. Such resonators underpin a wide range of applications, including dielectric and material-loss characterization [20, 8, 22], high-resolution spectroscopy [25, 11, 1, 31], cavity electromechanical [4, 9, 2], and precision frequency control [17]. They also serve as key elements in quantum devices, hybrid photonic-microwave systems, and precision tests of fundamental physics, where stability and phase coherence are paramount [27, 23].
The way in which two or more resonators are coupled governs the exchange of energy and phase information between them, determining the collective response of the system [16]. Controlled coupling enables a rich variety of interference effects, including linewidth narrowing, normal-mode splitting, mode hybridization, or sharp anti-resonances [5, 29, 3, 15]. By engineering the strength and phase of inter-resonator coupling, one can tailor the spectral response of the system for specific measurement or filtering purposes, as well as create conditions for coherent cancellation and enhanced sensitivity.
Coupled-resonator systems are conventionally realized through shared inductive or capacitive elements, or via aperture-mediated coupling mechanisms [25], each producing reactive interactions that modify the system’s collective resonance response. Resistive coupling, by contrast, has typically been regarded as a loss mechanism rather than a deliberate means of inter-resonator interaction. Some methods have been developed that invoke resistive coupling, including microwave filter networks [10], power-line communication systems [28], and cavity resonances coupling to line resonance due to a finite voltage standing wave ratio (VSWR) [30], though these are generally confined to planar or lumped-element implementations. To our knowledge, mutual resistive coupling between 3D microwave cavity resonators, realized here, has not been systematically investigated as a standalone coupling mechanism.
The motivation for this study arises from a three-cavity microwave system previously developed for tests of the scalar Electric Aharonov-Bohm effect [14, 13], in which two input cavities drove a third output cavity through thin copper foils many skin depths thick. When tuned, the system exhibited sharp anti-resonances with a strong dependence on the relative input phase.
[…]
Building upon these observations, we enhanced the readout sensitivity and developed an analytical model of the three-cavity system (Fig. 1) by introducing mutual resistance terms [30] to describe coupling through the foils. Each cavity is represented as a lumped LCR resonator, with resistive inter-cavity terms proportional to the dissipative power transferred across the interface. The model reproduces the observed interference and anti-resonant behaviour and is complemented by a first-principles derivation relating the foil’s electromagnetic properties to the measured coupling.
This work establishes a new coupling mechanism between volumetric microwave resonators, arising purely from resistive dissipation across a conductive interface. The framework provides both experimental and analytical foundations for resistive coupling in microwave resonators, enabling controlled anti-resonances and balance conditions useful for interferometric precision measurements and tests of fundamental physics.
II Equivalent Circuit Model
An equivalent circuit representation of the experimental system is shown in Fig. 2. Each of the three resonators in Fig. 1a is modelled as an LCR circuit characterised by its resonance frequency $f_{n}$, quality factor $Q_{n}$, and probe coupling coefficients $\beta_{n}$. The conductive foils separating the cavities introduce mutual resistive terms, $R_{13}$ and $R_{23}$, corresponding to the dissipative current flow between the input and output resonators.
Figure 2: The equivalent circuit diagram of the three-cavity resonator system coupled through thin metal foils. The three resonators are modeled by standard LCR circuits with capacitance $C_{n}$, inductance $L_{n}$, resistance $R_{n}$, and probe resistances $R_{0n}$ for resonator $n$. The mutual resistance terms $R_{13}$ and $R_{23}$, are shown coupling the resonators. The currents entering resonators 1 and 2, are modelled by current sources $I_{1}$ and $I_{2}$, with $I_{3}$ the current entering resonator 3 due to mutual contributions. The output current $I_{out}$ passes through the third resonator’s probe.
The transfer function, $T$, of the system is defined as
$T=\frac{I_{out}}{I_{in}},$ (1)
where phasors $I_{in}$ is the Thevenin equivalent input current at port 1, and $I_{out}$ is the current dissipated across the port 2 (resonator 3) output resistor. Under matched impedance conditions, this quantity is proportional to the forward transmission coefficient of the microwave circuit, $S_{21}$, and therefore represents the measured complex transmission response. The input current is split into $I_{1}$ and $I_{2}$, and a relative phase, $\phi$, and attenuation, $\alpha$, is introduced between the two new currents, related by
$I_{1}=\frac{1}{\sqrt{2}}I_{in}$ (2)
and
$I_{2}=\frac{1}{\sqrt{2}}\alpha e^{i\phi}I_{in}.$ (3)
Each of the parallel LCR circuits ($n=1,2,3$) has a self admittance approximated by $Y_{nn}\approx\frac{1}{Z_{nn}}$ in the small mutual coupling limit, where $Z_{nn}$ is the self impedance of cavity n. The admittance is given by
$Y_{nn}=\frac{1}{R_{0n}}+\frac{1}{R_{n}}+i\omega C_{n}+\frac{1}{i\omega L_{n}},$ (4)
with $R_{0n}$ and $R_{n}$ the probe and resonator resistances, respectively, capacitance $C_{n}$, inductance $L_{n}$, and angular frequency $\omega=2\pi f$. Near resonance, one can set $C_{n}\xrightarrow{}\frac{1}{\omega^{2}_{n}L_{n}}$, $L_{n}\xrightarrow{}\frac{R_{n}}{\omega_{n}Q_{n}}$ and $R_{n}\xrightarrow{}\beta_{n}R_{0n}$. Taking the power series and excluding terms quadratic and higher in $f$, the self-admittance becomes
$Y_{nn}=\frac{1}{\beta_{n}R_{0n}}(1+\beta_{n}+2iQ_{n}(\frac{f-f_{n}}{f_{n}})).$ (5)
The mutual admittances $Y_{13}=Y_{31}\approx\frac{-R_{13}}{R_{1}R_{3}}$ and $Y_{23}=Y_{32}\approx\frac{-R_{23}}{R_{2}R_{3}}$ (simplification justified in Section IV) couple resonators 1–3 and 2–3, respectively, while resonators 1 and 2 remain uncoupled with no mutual foils ($Y_{12}=Y_{21}=0$). It is here that we can introduce a normalised mutual coupling term $\Delta$ with reference to the mutual resistance
$\Delta_{n3}=\frac{R_{n3}}{R_{03}\sqrt{\beta_{n}\beta_{3}}}.$ (6)
Nodal analysis on the equivalent circuit in Fig. 2, gives three simultaneous equations:
$I_{1}$ $=Y_{11}V_{1}+Y_{13}V_{3},$ (7a)
$I_{2}$ $=Y_{22}V_{2}+Y_{23}V_{3},$ (7b)
$0$ $=Y_{31}V_{1}+Y_{32}V_{2}+Y_{33}V_{3},$ (7c)
for resonators 1, 2 and 3, respectively, with three unknown voltages $V_{1}$, $V_{2}$ and $V_{3}$ corresponding to each parallel node. Solving Eqs. (7a)–(7c) simultaneously, $T$ can be determined. Written in terms of microwave parameters, it becomes
$T=-\frac{1}{\sqrt{2}}\frac{f_{3}\sqrt{\beta_{3}}(f_{1}\sqrt{\beta_{1}}\gamma_{2}\Delta_{13}+e^{i\phi}\alpha f_{2}\sqrt{\beta_{2}}\gamma_{1}\Delta_{23})}{-\gamma_{1}\gamma_{2}\gamma_{3}+f_{1}\gamma_{2}f_{3}\Delta_{13}^{2}+\gamma_{1}f_{2}f_{3}\Delta_{23}^{2}},$ (8)
with the function
$\gamma_{n}=-2ifQ_{n}+f_{n}(-1-\beta_{n}+2iQ_{n}),$ (9)
for $n=1,2,3$.
Interesting phenomena arise when the currents through to the output resistor are closely balanced. To determine the balance condition, we set the attenuation factor $\alpha$ acting on $I_{2}$, such that $T=0$. Under the assumption that all three resonators are tuned to the same resonant frequency ($f_{1}=f_{2}=f_{3}$), and that the relative phase is matched, the required attenuation is,
$\alpha=\frac{\sqrt{\beta_{1}}(1+\beta_{2})\Delta_{13}}{\sqrt{\beta_{2}}(1+\beta_{1})\Delta_{23}}.$ (10)
Eq. 10 provides the symmetric-limit intuition; in practice small detunings shift the optimal balance point slightly, which is captured in the full model (Eq. 8) used for fitting.
III Fitting the Data
Three experimental spectra are plotted in Fig. 3 and fitted using the transfer function in Eq. (8) via standard nonlinear least-squares optimisation in Python. All measurements presented here were acquired with improved electromagnetic shielding and a lower-noise amplification chain, resulting in enhanced contrast and signal-to-noise ratio compared to [14, 13]. The resonance frequencies $f_{1}$ and $f_{2}$, and probe coupling coefficients $\beta_{1}$ and $\beta_{2}$, were determined independently from single-resonator characterisation and held fixed during the fit; all remaining parameters were extracted from the optimisation. The relative phase between the two input signals was varied from a case of destructive interference $(\phi=\pi)$ to constructive interference $(\phi=0)$. In the destructive case, a sharp cancellation (anti-resonance) feature emerges and the phase slope is enhanced compared to the near-resonant case.
Figure 3: Magnitude and phase versus frequency, with the best-fit transfer function from Eq. (8) plotted alongside the data. The corresponding fit parameters are listed in Table 1. Panels (a)-(c) illustrate the incremental approach towards resonance as the fitted relative phase $\phi$ is varied from $\pi$ to near $2\pi$ radians, respectively. The input power at port 1 was 20 dBm and the measured $S_{21}$ spectra are referenced to the input of the amplifier chain.
There is good agreement between the model and experiment, with the extracted parameters summarised in Table 1. The model reproduces this measured dependence on the relative input phase $\phi$: as $\phi$ is tuned towards the balanced condition, a sharp anti-resonance forms and the phase slope steepens.
Small system asymmetries (e.g., unequal resonance frequencies or $Q$-factors) create the ideal single anti-resonance when $\phi$ approaches $\pi$, as observed in Fig. 3a. Additional imbalance in $\alpha$ can lead to a splitting of this feature within the nearby phase region. In the symmetric limit, where all relevant parameters are equal, exact cancellation occurs at $(\phi=\pi)$ and broadband suppression is recovered across the plotted frequency span.
The central resonator frequency, $f_{3}$, was allowed to vary within a narrow range of approximately $\pm 30$ kHz during the fitting procedure, which is substantially smaller than the resonator bandwidth (of order MHz). This small freedom accounts for slight frequency drift observed during measurements, most likely arising from temperature or pressure fluctuations that becomes visible at the achieved low-noise measurement level. Other than $f_{3}$, only $\phi$ was allowed to vary over the range and all other parameters were kept constant. The model shows strong agreement with measurements across the full tuning range, as demonstrated in Fig. 4, which compares the measured response with the simulated transfer function as $\phi$ is swept over approximately $\pi$, consistent with the experimentally measured tuning range of $3.00$ rad at 6.1 GHz.
(a) Model
(b) Data
Figure 4: Comparison between the simulated transfer function (using the parameters in Table 1) for a near-$\pi$ span of $\phi$ (left) and the interpolated 98 measured spectra (right). The absolute phase reference is arbitrary; therefore a constant phase offset was applied for visual alignment between model and data. Both plots are shown for the experimentally determined $\phi$ range.
Table 1: System parameters used for the fits in Fig. 3.
Measured parameters
$f_{1}$ (GHz) $6.10883$
$f_{2}$ (GHz) $6.10886$
$\beta_{1}$ $0.81$
$\beta_{2}$ $1.26$
Fitted parameters
$f_{3}$ (GHz) $6.10883$
$Q_{1}$ $2.7\times 10^{3}$
$Q_{2}$ $3.1\times 10^{2}$
$Q_{3}$ $12.2\times 10^{3}$
$\Delta_{13}$ $(5.00\pm 0.01)\times 10^{-6}$
$\Delta_{23}$ $(4.10\pm 0.01)\times 10^{-6}$
$\alpha$ $1.23$
$\beta_{3}$ $0.93$
The three resonators exhibit comparable quality factors and are only weakly detuned, consistent with the near-symmetric behaviour expected for the balanced interference condition. The fitted mutual coupling coefficients are very small, on the order of $10^{-6}$, consistent with the weak interaction expected when the electromagnetic field couples across many skin depths of copper. The small difference between $\Delta_{13}$ and $\Delta_{23}$ likely arises from variations in surface field distributions, minor geometric asymmetries, or small differences in foil thickness. In Section IV, the physical interpretation of this mutual coupling is discussed and an expected value for the coupling rate based on mutual resistance is derived.
To directly measure the phase enhancement, an independent spectrum in Fig. 5 was acquired by finely tuning the balance parameters ($\phi$ and $\alpha$) to approach the destructive condition. This demonstrates a near order-of-magnitude increase in the local phase slope compared with the unbalanced configuration. This provides a direct, data-driven confirmation of the phase enhancement predicted by the model. Though both Fig. 5 and Fig. 3a show the destructive case, small differences in balance parameters can cause the polar plot to circle the origin and give the enhanced phase response in Fig. 5, rather than the inversion seen in Fig. 3a. This is similar to the difference between an under-coupled and over-coupled resonator.
Figure 5: Measured phase response near the balanced interference condition showing enhanced phase sensitivity. The local phase slope increases by nearly an order of magnitude relative to the near-resonant operating point shown in Fig. 3c, consistent with the behaviour expected near the antiresonance condition.
Due to ambient temperature and pressure fluctuations, maintaining the antiresonant condition is experimentally challenging, as the balanced interference point is extremely sensitive to small frequency drifts. Improved thermal and environmental stability, such as operation under vacuum or at cryogenic temperature, is expected to significantly improve the robustness of this operating regime and may be explored in future work.
Having established the phenomenology of the enhancement, we next consider the physical origin of the weak mutual coupling responsible for this behaviour.
IV Coupling Across a Metallic Foil
IV.1 Derivation
To verify the physical significance of the mutual coupling term $\Delta_{n3}$, and hence the mutual resistance $R_{n3}$, we derive its value from first principles and compare it with the value obtained from model fitting. First, consider the electric field induced at either side of a thin sheet, $\tilde{E}_{n}$ and $\tilde{E}_{3}$, for foil sides n (where n is now either 1 or 2, representing the side of the associated input resonator) and 3, respectively:
$\tilde{E}_{n}=Z_{s}\tilde{J}_{sn}+Z_{t}\tilde{J}_{s3},\quad$ (11)
$\tilde{E}_{3}=Z_{s}\tilde{J}_{s3}+Z_{t}\tilde{J}_{sn},\quad$ (12)
with surface currents $\tilde{J}_{sn}$ and $\tilde{J}_{s3}$ at sides n and 3, respectively, self impedance $Z_{s}$, at the two sides, and $Z_{t}$ is the transfer impedance. The expression for $Z_{t}$ is obtained from a transition boundary condition [7], which models a geometrically thin, but not electrically thin, sheet. To our knowledge, such a derivation is not readily available in the literature, and is thus provided in Appendix A. Here, the self and transfer impedances are defined as
$Z_{s}=-\frac{i\omega\mu}{k}\frac{1}{tan(kd)},\quad$ (13)
$Z_{t}=-\frac{i\omega\mu}{k}\frac{1}{sin(kd)},\quad$ (14)
with propagation distance (or foil thickness) d. The plane wave propagation constant $k$, is given by $k=\omega\sqrt{(\epsilon+(\frac{\sigma}{i\omega}))\mu}$, with angular frequency $\omega$, conductivity $\sigma$, and the media’s permittivity and permeability $\epsilon$ and $\mu$, respectively. Only the dissipative part contributes to mutual power transfer; therefore the real part of $Z_{t}$ is retained.
The dissipated power PdPd, is generally
$Pd=\int Re(\tilde{E}.\tilde{J}_{s}^{*})dS,\quad$ (15)
which after substituting Eqs. 11 and 12 becomes
$Pd=\int(Z_{s}|\tilde{J}_{sn}|^{2}+Z_{s}|\tilde{J}_{s3}|^{2}+2Re(Z_{t})Re(\tilde{J}_{sn}\tilde{J}_{s3}^{*}))dS,\quad$ (16)
where the final term is the mutual dissipated power
$Pd_{n3}=2Re(Z_{t})\int Re(\tilde{J}_{sn}.\tilde{J}_{s3}^{*})dS.\quad$ (17)
This can also be written as
$Pd_{n3}=Y_{n3}Re(V_{n}V_{3}^{*}),\quad$ (18)
with the voltages $V_{n}$ and $V_{3}$ at sides n and 3, respectively, and in terms of the mutual admittance $Y_{n3}$. Combining Eq 17 and 18, we can solve for the mutual admittance
$Y_{n3}=2Re(Z_{t})\frac{\int Re(\tilde{J}_{sn}.\tilde{J}_{s3}^{*})dS}{Re(V_{n}V_{3}^{*})}.\quad$ (19)
Substituting the stored energy of the microwave resonator $U=\frac{C|V|^{2}}{2}$, with resonator capacitance $C$, and noting $V=|V|e^{i\phi}$ with phase $\phi$, we can write the mutual admittance as
$Y_{n3}=Re(Z_{t})\sqrt{\frac{C_{n}C_{3}}{U_{n}U_{3}}}\frac{\int Re(\tilde{J}_{sn}.\tilde{J}_{s3}^{*})dS}{Re(e^{i(\phi_{n}-\phi_{3})})}.\quad$ (20)
The subscripts of $C$, $U$, and $\phi$ refer to the resonator associated to that side. $\tilde{J}_{sn}$ and $\tilde{J}_{s3}$ can be related to the magnetic fields at their respective surfaces, giving
$Y_{n3}=Re(Z_{t})\sqrt{\frac{C_{n}C_{3}}{U_{n}U_{3}}}\frac{\int Re((\vec{n}_{n}\times\vec{H}_{tn}).(\vec{n}_{3}\times\vec{H}_{t3})^{*})dS}{Re(e^{i(\phi_{n}-\phi_{3})})},\quad$ (21)
with $\vec{n}_{n}$ and $\vec{n}_{3}$ unit vectors normal to their respective side in the $\pm z$ directions. Near both sides of the foil, the tangential magnetic field of a cylindrical resonator operating in the TM010 is
$\vec{H}_{t}=-E_{0}\begin{pmatrix}0\\ i\omega\epsilon\frac{a}{\chi_{01}}J_{0}^{\prime}(\frac{\chi_{01}r}{a})\\ 0\end{pmatrix}e^{i\phi},\quad$ (22)
at a radius $r$, with electric field normalisation constant $E_{0}$ and cavity radius $a$ for the respective resonator. $\vec{H}_{tn}$ and $\vec{H}_{t3}$ involve the microwave experimental parameters for their respective sides. $J_{m}$ is the Bessel function of the m-th kind where the primed $J_{m}^{\prime}$ is the derivative, and $\chi_{mu}$ is the u-th root of the Bessel functions of the m-th kind. This mode has capacitance
$C=\frac{\pi\epsilon_{0}a^{2}J_{1}(\chi_{01})^{2}}{h}\quad$ (23)
and stored energy
$U=\frac{1}{2}\pi\epsilon_{0}a^{2}hE_{0}^{2}J_{1}(\chi_{01})^{2},\quad$ (24)
with cavity height $h$ and permittivity of free space $\epsilon_{0}$. Since, in this case, the metallic foil couples the ends of two cylindrical resonators of different radii, the integral is computed over the smaller radius. Substituting Eqs. 22, 23 and 24 into 21 and using the integral
$\int J_{0}^{\prime}(\frac{\chi_{01}r}{a_{n}})J_{0}^{\prime}(\frac{\chi_{01}r}{a_{3}})rdr=-\frac{r(a_{n}J_{0}(\frac{r\chi_{01}}{a_{3}})J_{1}(\frac{r\chi_{01}}{a_{n}}))}{(\frac{a_{n}}{a_{3}}-\frac{a_{3}}{a_{n}})\chi_{01}},$ (25)
the admittance becomes
$Y_{n3}=-4\pi\omega^{2}\epsilon^{2}\frac{Re(Z_{t})a_{n}^{4}a_{3}J_{0}(\frac{a_{n}\chi_{01}}{a_{3}})J_{1}(\chi_{01})}{\chi_{01}^{3}h_{n}h_{3}(a_{n}^{2}-a_{3}^{2})}.$ (26)
The subscript of the $a$ and $h$ terms refer the radii and height of the respective cavity. Solving the standard admittance matrix we can calculate the mutual resistance as
$R_{n3}=\frac{1\pm\sqrt{1+4Y_{n3}^{2}R_{n}R_{3}}}{2Y_{n3}}\approx-Y_{n3}R_{n}R_{3},$ (27)
with the approximation valid for $4Y_{n3}^{2}R_{n}R_{3}<<1$ in the negative root, indeed the case for the parameters shown in Table 2 and with frequencies around 66 GHz. $R_{n}$ and $R_{3}$ are the parallel resistance of the respective resonator given by
$R_{k}=\frac{h_{k}^{2}\mu_{0}}{2\pi a_{k}(a_{k}+h_{k})R_{surface}\epsilon_{0}J_{1}(\chi_{01})^{2}}.$ (28)
$\mu_{0}$ is the permeability of free space, and $R_{surface}$ is the surface resistance $R_{surface}=\sqrt{\frac{\omega\mu_{0}\rho}{2}}$ with resistivity $\rho$. Thus, Eq. 27 combined with Eq’s. 26 and 28 determines the mutual resistance across the foil connecting two resonators.
IV.2 Calculated Values
The physical parameters of the experiment are shown in Table 2. The foil thickness was verified using a micrometer. The resonator parameters in conjunction with a few physical constants can be plugged in to Eq. 27 to obtain an expected mutual resistance and thus coupling.
Table 2: Physical constants and parameters of the experiment.
Parameter Value
$a_{n}$ $(18.00\pm 0.005)\times 10^{-3}\ \text{m}$
$a_{2}$ $(18.75\pm 0.005)\times 10^{-3}\ \text{m}$
$h_{n}$ $(35.00\pm 0.01)\times 10^{-3}\ \text{m}$
$h_{2}$ $(38.64\pm 0.01)\times 10^{-3}\ \text{m}$
$d$ $(9\pm 1)\times 10^{-6}\ \text{m}$
$\varepsilon_{0}$ $8.854\times 10^{-12}\ \text{F/m}$
$\mu_{0}$ $4\pi\times 10^{-7}\ \text{H/m}$
$\rho_{\text{Cu}}$ $1.68\times 10^{-8}\ \Omega\,\text{m}$
$\sigma_{\text{Cu}}$ $5.96\times 10^{7}\ \text{S/m}$
The mutual resistance $R_{n3}$ is calculated to lie in the range $R_{n3}\approx(6\text{--}300)\,\Omega$. The negative root of Eq. 27 is taken as the physical solution; the sign not meaningful, so the magnitude is reported. This extremely small value is intuitively consistent with the parallel-circuit picture: only small current flows across the foil, so the resulting coupling is weak, when the foil thickness spans multiple skin depths. Recalling Eq. 6, and expressing the resonator resistance as $R_{n}=R_{0n}\beta_{n}$, the mutual coupling is calculated to lie in the range $\Delta_{n3}\approx(1\text{--}48)\times 10^{-6}$ when the estimated foil thickness of $(9\pm 1)\,\mu\mathrm{m}$ is considered. Comparing this with the experimentally fitted values of $\Delta_{13}=(5.00\pm 0.01)\times 10^{-6}$ and $\Delta_{23}=(4.10\pm 0.01)\times 10^{-6}$ shows good agreement within the expected uncertainty. Given the strong sensitivity of the coupling to foil thickness and other geometric parameters, agreement at this level constitutes successful verification of the model.
The remaining variation between the expected and measured coupling values likely arises from uncertainties in highly sensitive parameters, such as the foil thickness, where a difference of a few micrometres can result in an order-of-magnitude change in the coupling. The foil thickness is considered to be uniform for the purposes of the calculation. More precise measurements of foil thickness would further reduce this uncertainty. Additionally, the resonator parameters such as conductivity may vary due to oxidation and other cavity differences. Small variations in the resonator mode structures will also lead to variations in the calculated cavity resistances.
V Conclusion
We develop an equivalent circuit model to describe the coupling between three microwave cavity resonators interconnected by thin metallic foils. By introducing mutual resistance terms, the model accurately reproduces the experimentally observed interference phenomena and anti-resonant behaviour. The derived coupling values agree with those extracted from the fitted transfer function, confirming the physical validity of resistive inter-cavity coupling.
This study demonstrates a previously unexplored coupling mechanism between volumetric microwave resonators, arising purely from dissipative interaction across a continuous conductive interface. This new approach enables precise control of weak coupling and the formation of sharp anti-resonances, in contrast to traditional aperture-based and other established coupling methods.
The demonstrated model offers a robust foundation for the design of new resonator architectures in precision microwave experiments. Such resistive coupling mechanisms may be exploited in fundamental physics tests sensitive to phase and potential differences, including studies of the scalar Electric Aharonov-Bohm effect and related dark-sector interactions. Future work may involve testing different foil and mode geometries for enhanced, rather than destructive, coupling. Additionally, this phenomenon can be investigated at cryogenic temperatures. Collectively, these results provide a rigorous theoretical and experimental basis for employing metallic foil mutual resistance as a controllable coupling mechanism in multi-resonator microwave systems.
Acknowledgements
This work was funded by the ARC Centre of Excellence for Dark Matter Particle Physics, CE200100008, and by the Defence Science Centre, an initiative of the State Government of Western Australia.
We thank Eugene N. Ivanov for valuable assistance in experimental preparation.
Appendix A Transition Boundary Condition Derivation
Across a thin material layer, the electric fields tangential to the surface on side 1 and side 2, $E_{t1}$ and $E_{t2}$, respectively, are given by
$\tilde{E}_{t1}$ $=$ $Z_{s}\tilde{J}_{s1}+Z_{t}\tilde{J}_{s2},\quad$ (29)
$\tilde{E}_{t2}$ $=$ $Z_{t}\tilde{J}_{s1}+Z_{s}\tilde{J}_{s2}.\quad$ (30)
$Z_{s}$ and $Z_{t}$ are the self and transfer impedances of the material layer, and $J_{s1}$ and $J_{s2}$ the surface current densities on the two sides. This comes from the fact that the boundary condition due to a large conductivity allows us to generally write, with little error [6], $\tilde{E}_{t}=Z_{m}J_{s}$, where $Z_{m}$ is the surface resistance. Solving for the surface currents in Eqs. 29 and  30, we arrive at[7]
$\tilde{J}_{s1}=\frac{Z_{s}\tilde{E}_{t1}-Z_{t}\tilde{E}_{t2}}{Z_{s}^{2}-Z_{t}^{2}},\quad$ (31)
$\tilde{J}_{s2}=\frac{Z_{s}\tilde{E}_{t2}-Z_{t}\tilde{E}_{t1}}{Z_{s}^{2}-Z_{t}^{2}}.\quad$ (32)
Following the standard transition matrices one has
$\begin{bmatrix}V_{1}\\ I_{1}\end{bmatrix}=\begin{bmatrix}\cosh(\gamma d)&Z_{0}\sinh(\gamma d)\\ \frac{1}{Z_{0}}\sinh(\gamma d)&\cosh(\gamma d)\end{bmatrix}\begin{bmatrix}V_{2}\\ I_{2}\end{bmatrix}$ (33)
and
$\begin{bmatrix}V_{1}\\ V_{2}\end{bmatrix}=\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}\end{bmatrix}\begin{bmatrix}I_{1}\\ I_{2}\end{bmatrix}$, (34)
for voltages and currents $V_{n}$ and $I_{n}$, respectively. Here, $d$ is the propagation distance, and $\gamma=i\omega\sqrt{(\epsilon+\frac{\sigma}{\omega i})\mu}$ is the propagation constant (wave number) determined by the material permittivity, conductivity, and permeability. One can then solve for $Z_{s}$ and $Z_{t}$ by defining $Z_{s}=Z_{11}$ and $Z_{t}=-Z_{12}$ arriving at
$Z_{s}=-iZ_{0}\frac{1}{tan(kd)},\quad$ (35)
$Z_{t}=-iZ_{0}\frac{1}{sin(kd)},\quad$ (36)
setting $\gamma=ik$. Finally, the characteristic impedance is given by [6]
$Z_{0}=\sqrt{\frac{\mu}{(\epsilon+\frac{\sigma}{i\omega})}}.$ (37)
And thus, we have derived the transition boundary condition [7].